OpenStudy (anonymous):

find the critical numbers f(x)=-x^4+4x^3+5

Parth (parthkohli):

As they always say: find \(f'(x)\) and equate that to \(0\).

Parth (parthkohli):

@chochko What would \(f'(x)\) be here?

OpenStudy (anonymous):

-4x^3+12x^2

OpenStudy (anonymous):

ANd then?

Parth (parthkohli):

Well that's correct - now equate it to zero.

Parth (parthkohli):

Solve this equation:\[\rm 4x^3 + 12x^2 = 0\]

OpenStudy (anonymous):

-4x^3+12x^2=0

Parth (parthkohli):

Yeah, I meant -4x^3 + 12x^2 = 0

OpenStudy (anonymous):

sry my computer is really slow. Would you factor out a 4?

OpenStudy (anonymous):

Yu.. 4x^2 ( x-3) = 0 4x^2 = 0 x = 0 x-3 = 0 x = 3

Parth (parthkohli):

Mine too, and you can factor 4x^2 out.

Parth (parthkohli):

Yep, and what Yahoo! said are the critical points. ;)

OpenStudy (anonymous):

thank you but im not done . that is just one part..can you help yahoo or ParthKohli?

Parth (parthkohli):

Sure.

OpenStudy (mimi_x3):

the next part is to differentiate it again :P to determine its nature

Parth (parthkohli):

You guys must be in the same school.

Parth (parthkohli):

lol

OpenStudy (anonymous):

using the second derivitive test which is -12x^2+24x=0 set to 0 to determine all relative extrema, indicating the x and Y values and wheteher is a max or a min. Help please

OpenStudy (mimi_x3):

lol @ParthKohli: the next step is common sense. @chochko: with the differentiataed function sub the \(x\) that you found in the first step if its a negative then its concave down => max. if its concave up then => min

Parth (parthkohli):

Sorry, was on another question, and yeah - that's the way I was taught too.

OpenStudy (anonymous):

so in -12x^2+24x Sub the x for the critical numbers and the result is the relative extrema? is that the Max and Min as well?

OpenStudy (mimi_x3):

max or min concave up => min concave down => max

OpenStudy (anonymous):

so which part is the relative extrema? i really suck at math btw?

OpenStudy (mimi_x3):

well i googled; relative extrema and max/ min extrema is the same thing

OpenStudy (anonymous):

f''(0)=0 min f''(3)=-36 Max Is this correct?

OpenStudy (mimi_x3):

i dont know; i dont have a calculator with me but if its you have to check if it has a horizontal point of inflexion

OpenStudy (anonymous):

Ok, so the intervals are the critical points right? where are they increasing and decreasing?

OpenStudy (mimi_x3):

critical points are the stationary points on the curve since \(f(x)=0\) you differentiate it twice to determine if its increasing or decreasing increasing \(f''(x) >0\)

OpenStudy (mimi_x3):

i think i made a mistake; increasing \(f'(x)>0\)

OpenStudy (mimi_x3):

when the function is decreasing \(f'(x)<0\)

OpenStudy (mimi_x3):

i mean you differetiate it once; to see if its increasing or decreasing sorry; im sick

OpenStudy (anonymous):

so using the f'(x)=0 and plug in 0 and 3 for x to find where f(x) is increasing or decreasing?

OpenStudy (mimi_x3):

oh no no im getting tired @ParthKohli: may like to explain :)

OpenStudy (anonymous):

ok thanks a lot

OpenStudy (anonymous):

@ParthKohli are you able to help me finish this one

OpenStudy (mimi_x3):

I think I should explain; since Parth is hiding :P \(f'(x) = 0\) you are finding the \(x\) values that is the STATIONARY point. \(f'(x) >0\) is the \(x\) values were the function is INCREASING \(f'(x)<0\) is where the function is DECREASING. The second derivative is the deirivative of the derivative. The curve \(y=f(x)\) is concave up when \(f''(x)>0\) and concve down when \(f''(x)<0\)

Parth (parthkohli):

Sry, back.

Parth (parthkohli):

Didn't notice the tag right there. :P

OpenStudy (anonymous):

Increasing on ( -inf, 0) (0,3) then decreasing (3, inf) idk if its correct?

Parth (parthkohli):

Well your \(\rm f'(x)\) function is \(\rm -4x^2 +12x \) and the function is stationary at x = 3,0.

Parth (parthkohli):

What is the question again?

OpenStudy (anonymous):

State the intervas for which f(x) is increasing or decreasing So i tested the values -1, 1, 4 And got F'(-1)=16 inc f'(2)=8 inc f'(4)=-64 dec. idk if its correct?

Parth (parthkohli):

That's right.

Parth (parthkohli):

So it starts decreasing after 3.

OpenStudy (anonymous):

(-inf, 0) (0,3) increasing then (3, inf) decreasing Are the intervals right?

Parth (parthkohli):

Let's test f'(-5). f'(x) = -4x^2 + 12x f'(-5) = -4(-5)^2 + 12(-5) = -4(25) - 60 = -100 - 60 = -160

Parth (parthkohli):

It's decreasing at f(-5).

Parth (parthkohli):

The best way is to ask Wolfram for the graph. :)

OpenStudy (anonymous):

huh? whos that?

Parth (parthkohli):

http://wolframalpha.com

Parth (parthkohli):

Enter your function and see its graph.

OpenStudy (anonymous):

ok can you help me find the possible inflection points

OpenStudy (anonymous):

@ParthKohli can you help just a lil bit left

Parth (parthkohli):

Sorry, back.

Parth (parthkohli):

Damn... I've disabled the tag notifs.